# Distributionally Robust Markov Games with Average Reward

### v1

@@ 6 Algorithmic Solutions
We then develop an algorithmic solution for average-reward DR-MGs. Notably, most of the existing algorithms for non-robust average-reward games are developed for two-player zero-sum MGs Filar and Vrieze (2012), which cannot be extended to the robust setting due to the existence of the additional hidden environment player Zhang et al. (2020). For example, standard policy iteration can not guarantee a monotonous performance due to the robustness aspect. We propose a value iteration based algorithm in Algorithm 1, named robust Nash-Iteration.
@@
Our algorithm maintains an estimation, $h_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, achieving such a Nash Equilibrium can be PPAD-complete Deng et al. (2023). Moreover, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies Li (2003); Hu and Wellman (2003); Littman et al. (2001); Bowling (2001, 2000). This assumption is not merely a technical convenience; it is the conceptual linchpin that stabilizes the learning dynamics of the system. By placing constraints on which Nash Equilibrium the oracle returns, we ensure that the sequence of generated policies and value functions behaves predictably, guiding the algorithm steadily toward a solution instead of allowing it to wander chaotically or diverge. We then derive the convergence of our algorithm.
@@
With a specific equilibrium selection oracle NE (see Assumption D.1 in the Appendix), the robust Nash-Iteration algorithm converge to a Nash Equilibrium of the average-reward distributionally robust game.
@@
7 Connections with Discounted Nash Equilibrium
In this section, we study the connection between distributionally robust games with discounted and average rewards. Our motivation is to utilize the extensively studied and mathematical elegance of the discounted setting to solve the more challenging average-reward setting. Recall that a policy $\pi^{*}_{\gamma}$ is a Nash Equilibrium of the $\gamma$-discounted DR-MG if
$V^{\pi^{*}}_{\mathcal{P},i}\geq V^{(\pi^{*}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (9)
where $V^{\pi^{*}}_{\mathcal{P},i}$ is the $\gamma$-discounted robust value function. A policy $\pi^{*}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{*}}_{\mathcal{P},i}+\epsilon\geq g^{(\pi^{*}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (10)
We then derive the following results.
@@ 7 Connections with Discounted Nash Equilibrium
The importance of this theorem lies in its computational implications. The analysis of $\gamma$-discounted games is often more tractable Kardeş, Ordóñez, and Hall (2011), benefiting from properties like the Banach fixed-point theorem which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as value iteration and policy iteration, are naturally formulated for the discounted setting. Our results hence provide a rigorous justification for using these well-understood tools to solve the more challenging average reward problem.
Remark 7.2.
@@
Define the robust diameter Auer, Jaksch, and Ortner (2008) of the distributionally robust Markov game by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
The result hence specifies a reduction factor whose Nash Equilibrium is an approximated equilibrium under the average reward criterion. Thus, we can apply existing algorithms for discounted DR-MGs Zhang et al. (2020) to approximately solve the average reward ones.
@@ 8.1 Long-Term Performance
To empirically validate our approach, we perform a series of experiments designed to compare the performance of our robust Nash-Iteration algorithm with average reward against the standard discounted-reward robust Nash-Iteration Zhang et al. (2020); Shi et al. (2024b) under DR-MGs. We aim to show that our algorithm will yield a joint policy that has better long-term performance than those derived from the discounted criterion.
@@
(1). The optimal average-reward policy, $\pi_{avg}$, is computed using the average-reward Robust Nash-Iteration algorithm. We evaluate the worst-case average reward of Player 1 under the policy, using the robust average reward single-agent RL method in Wang et al. (2023a).
@@ 8.2 Robustness
We then test the robustness of our method by comparing it with the non-robust average-reward Nash-Iteration algorithm Li (2003), to test their ability to handle model uncertainty. Using the same environment, we evaluate both learners against a worst-case adversary throughout training to measure their inherent robustness. The results are presented in Figure 2. The policy learned by the robust algorithm consistently and monotonically improves its worst-case performance, rapidly converging to a high, stable value. This is because the algorithm directly optimizes for resilience against the adversary at each step. In contrast, the policy learned by the non-robust algorithm exhibits lower performance when evaluated against a worst-case adversary. Ultimately, it does not have the ability to reliably account for perturbations in the underlying environment, as opposed to its robust counterpart, which aligns with our theoretical results.
@@ 9 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We established the well-posedness of the robust average-reward Bellman equation, proved the existence of a stationary Nash Equilibrium, and proposed the Robust Nash-Iteration algorithm as the first provably convergent method for this game class. We also connected average-reward equilibria to their discounted counterparts, and our numerical experiments validated the superior long-term performance and stability of our approach. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
@@ A.1 Discounted Setting
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator Nilim and El Ghaoui (2004); Iyengar (2005); Puterman (1994):
@@
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{p\in\mathcal{P}^{a}_{s}}p^{\top}V$ is the support function of $V$ on $\mathcal{P}^{a}_{s}$. Based on the contraction of $\mathbf{T}_{\pi}$, robust dynamic programming approaches, e.g., robust value iteration, can be designed Nilim and El Ghaoui (2004); Iyengar (2005).
@@
$\pi^{*}=\arg\max_{\pi}V^{\pi}_{\mathcal{P},\gamma}(s),$ (13)
and it is shown Iyengar (2005) that the optimal robust value function $V^{\pi^{*}}_{\mathcal{P},\gamma}(s)$ satisfies the optimal robust Bellman equation
@@ A.2 Average Reward Setting
We can view equation (16) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by Puterman (1994), thus it holds the following non-robust Bellman equation:
@@
It is shown that under the unichain assumption, the robust average reward is independent from the initial state: $g_{\mathcal{P}}^{\pi}(s_{1})=g_{\mathcal{P}}^{\pi}(s_{2})$, for any $s_{1},s_{2}$. The robust Bellman equation (for a policy $\pi$) and the robust optimal Bellman equation for the average reward setting are derived in Wang et al. (2023a):
@@
It is shown Wang et al. (2023b) that if the equation (19) ((20)) has a solution $(g,V)$, then the solution $g$ is the robust average reward $g^{\pi}_{\mathcal{P}}$ (the optimal robust average reward $g^{*}_{\mathcal{P}}=\max_{\pi}g^{\pi}_{\mathcal{P}}$). However, whether these two equations are solvable is not studied but directly assumed.
Appendix B Proof of Section 4
Theorem B.1.
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 4.1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the robust Bellman equation:
@@ Appendix B Proof of Section 4
We denote $s^{*}\triangleq\arg\min_{s\in\mathcal{S}}x(s)$, then it holds that
$x(s^{*})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s).$ (30)
By our assumption, for any $s^{\prime}\in\mathcal{S}$, there exists some $k$ such that $((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}>0$. Then it holds that
$x(s^{*})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s)\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s^{*})=x(s^{*}),$ (31)
@@
$x(s^{*})=((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime})+(1-((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}})x(s^{*}).$ (32)
@@
$((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{*})\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime}),$ (33)
which implies $x(s^{*})=x(s^{\prime}),\forall s^{\prime}\in\mathcal{S}$, and thus $x=me$ for some constant $m$, i.e., $V^{\pi}_{\mathsf{P}_{w}}=me+V^{\pi}_{\mathsf{P}_{V}}$.
@@
Denote the optimal robust policy by $\pi^{*}$, and the corresponding worst-case kernel by $\mathsf{P}^{*}$. By Theorem 4.3, $(g^{\pi^{*}}_{\mathcal{P}},h^{\pi^{*}}_{\mathsf{P}^{*}})$ is the solution to the robust Bellman equation of $\pi^{*}$:
$h^{\pi^{*}}_{\mathsf{P}^{*}}(s)=r^{\pi^{*}}(s)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{\pi^{*}}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})=\sum_{a}\pi^{*}(a|s)\left(r(s,a)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\right).$ (41)
We now set $\pi^{\prime}(s)=\arg\max_{a}\{r(s,a)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\},\forall s$, then we have that
$r^{\pi^{\prime}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})=g^{\pi^{\prime}}_{\mathcal{P}}\leq g^{\pi^{*}}_{\mathcal{P}}$ $=\sum_{a}\pi^{*}(a|s)\left(r(s,a)-h^{\pi^{*}}_{\mathsf{P}^{*}}(s)+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\right)$
$\leq r^{\pi^{\prime}}(s)-h^{\pi^{*}}_{\mathsf{P}^{*}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}}),$ (42)
@@
$h^{\pi^{*}}_{\mathsf{P}^{*}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)\leq\sigma^{\pi^{\prime}}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})-\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})\leq\mathsf{P}^{\prime}(h^{\pi^{*}}_{\mathsf{P}^{*}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}).$ (43)
Now denote $x=h^{\pi^{*}}_{\mathsf{P}^{*}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}$, it holds that

### v2

@@ 6 Algorithmic Solutions
We then develop an algorithmic solution for average-reward DR-MGs. Notably, finding stationary NE is significantly challenging in general-sum games, even under non-robust settings (hu2003nash; li2007reinforcement; li2003learning; zinkevich2005cyclic), and existing convergence results generally require additional structural assumptions on the underlying games. In this section, we design a value iteration based algorithm in Algorithm 1, named robust Nash-Iteration, to find a NE. The convergence guarantee of our algorithm, similarly to previous ones for general-sum games, is only obtained under some additional assumptions. Our algorithm is designed for irreducible DR-MGs, as learning under weakly communicating settings is generally challenging and remains open (sahabandu2024rl).
@@
Our algorithm maintains an estimation, $h_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). This assumption is not merely a technical convenience; it is the conceptual linchpin that stabilizes the learning dynamics of the system. By placing constraints on which Nash Equilibrium the oracle returns, we ensure that the sequence of generated policies and value functions behaves predictably, guiding the algorithm steadily toward a solution instead of allowing it to wander chaotically or diverge.
@@
Under ˜1, and with a specific equilibrium selection oracle NE (see Assumption 3 in Appendix), the robust Nash-Iteration algorithm converges to a Nash Equilibrium.
@@
7 Connections with Discounted DR-NE
In this section, we study the connection between distributionally robust games with discounted and average rewards. Our motivation is to utilize the extensively studied and mathematical elegance of the discounted setting to solve the more challenging average-reward setting. Recall that a policy $\pi^{\star}_{\gamma}$ is a Nash Equilibrium of the $\gamma$-discounted DR-MG if
$V^{\pi^{\star}}_{\mathcal{P},i}\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (9)
where $V^{\pi^{\star}}_{\mathcal{P},i}$ is the $\gamma$-discounted robust value function. A policy $\pi^{\star}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{\star}}_{\mathcal{P},i}+\epsilon\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (10)
We then derive the following results for irreducible DR-MGs (˜1).
@@ 7 Connections with Discounted DR-NE
The importance of this theorem lies in its computational implications. The analysis of $\gamma$-discounted games is often more tractable (kardecs2011discounted), benefiting from properties like the Banach fixed-point theorem which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as value iteration and policy iteration, are naturally formulated for the discounted setting. Our results hence provide a rigorous justification for using these well-understood tools to solve the more challenging average reward problem.
Remark 3.
@@
Define the robust diameter (auer2008near) of the distributionally robust Markov game by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
The result hence specifies a reduction factor whose Nash Equilibrium is an approximated equilibrium under the average reward criterion. Thus, we can apply existing algorithms for discounted DR-MGs (zhang2020robust) to approximately solve the average reward ones.
@@ 8.1 Long-Term Performance
To empirically validate our approach, we perform a series of experiments designed to compare the performance of our robust Nash-Iteration algorithm with average reward against the standard discounted-reward robust Nash-Iteration (zhang2020robust; shi2024sample) under DR-MGs. We aim to show that our algorithm will yield a joint policy that has better long-term performance than those derived from the discounted criterion.
@@
(1). The optimal average-reward policy, $\pi_{avg}$, is computed using the average-reward Robust Nash-Iteration algorithm. We evaluate the worst-case average reward of Player 1 under the policy, using the robust average reward single-agent RL method in (wang2023robust).
@@ 8.2 Robustness
We then test the robustness of our method by comparing it with the non-robust average-reward Nash-Iteration algorithm (li2003learning), to test their ability to handle model uncertainty. Using the same environment, we evaluate both learners against a worst-case adversary throughout training to measure their inherent robustness. The results are presented in Figure 2. The policy learned by the robust algorithm consistently and monotonically improves its worst-case performance, rapidly converging to a high, stable value. This is because the algorithm directly optimizes for resilience against the adversary at each step. In contrast, the policy learned by the non-robust algorithm exhibits lower performance when evaluated against a worst-case adversary. Ultimately, it does not have the ability to reliably account for perturbations in the underlying environment, as opposed to its robust counterpart, which aligns with our theoretical results.
@@ 9 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We established the well-posedness of the robust average-reward Bellman equation, proved the existence of a stationary Nash Equilibrium under both irreducible and weak communicating settings, and proposed the Robust Nash-Iteration algorithm as the first provably convergent method for this game. We also connected average-reward equilibria to their discounted counterparts, and our numerical experiments validated the superior long-term performance and stability of our approach. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
@@ A.1 Discounted Setting
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator (nilim2004robustness; iyengar2005robust; puterman1994markov):
@@
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{p\in\mathcal{P}^{a}_{s}}p^{\top}V$ is the support function of $V$ on $\mathcal{P}^{a}_{s}$. Based on the contraction of $\mathbf{T}_{\pi}$, robust dynamic programming approaches, e.g., robust value iteration, can be designed (nilim2004robustness; iyengar2005robust).
@@
$\pi^{\star}=\arg\max_{\pi}V^{\pi}_{\mathcal{P},\gamma}(s),$ (13)
and it is shown (iyengar2005robust) that the optimal robust value function $V^{\pi^{\star}}_{\mathcal{P},\gamma}(s)$ satisfies the optimal robust Bellman equation
@@ A.2 Average Reward Setting
We can view equation (16) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by (puterman1994markov), thus it holds the following non-robust Bellman equation:
@@
It is shown that under the unichain assumption, the robust average reward is independent from the initial state: $g_{\mathcal{P}}^{\pi}(s_{1})=g_{\mathcal{P}}^{\pi}(s_{2})$, for any $s_{1},s_{2}$. The robust Bellman equation (for a policy $\pi$) and the robust optimal Bellman equation for the average reward setting are derived in (wang2023robust):
@@
It is shown (wang2023model) that if the equation (19) ((20)) has a solution $(g,V)$, then the solution $g$ is the robust average reward $g^{\pi}_{\mathcal{P}}$ (the optimal robust average reward $g^{*}_{\mathcal{P}}=\max_{\pi}g^{\pi}_{\mathcal{P}}$). However, whether these two equations are solvable is not studied but directly assumed.
Appendix B Proof of Lemma˜3.1
We derive the result by constructing an example.
Fix $\varepsilon\in(0,1)$. Consider a two-player zero-sum distributionally robust Markov game
$(\mathcal{S},\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{P},r_{1},r_{2}),$
with state space $\mathcal{S}=\{s,\,G,\,B\}$, action sets $\mathcal{A}_{1}=\{1,2\}$ and $\mathcal{A}_{2}=\{1,2\}$, and rewards
$r_{1}(G,a_{1},a_{2})=1,\qquad r_{1}(B,a_{1},a_{2})=0,\qquad r_{1}(s,a_{1},a_{2})=0,\qquad r_{2}=-r_{1}.$
The ambiguity set consists of two kernels $p^{(1)}$ and $p^{(2)}$:
$p^{(1)}(\,\cdot\,|s,(1,1))$ $=\delta_{G},$ $p^{(1)}(\,\cdot\,|s,(2,2))$ $=\varepsilon\,\delta_{G}+(1-\varepsilon)\,\delta_{B},$
$p^{(2)}(\,\cdot\,|s,(1,1))$ $=\varepsilon\,\delta_{G}+(1-\varepsilon)\,\delta_{B},$ $p^{(2)}(\,\cdot\,|s,(2,2))$ $=\delta_{G},$
$p^{(1)}(\,\cdot\,|s,(1,2))$ $=\delta_{B},$ $p^{(1)}(\,\cdot\,|s,(2,1))$ $=\delta_{B},$
$p^{(2)}(\,\cdot\,|s,(1,2))$ $=\delta_{B},$ $p^{(2)}(\,\cdot\,|s,(2,1))$ $=\delta_{B},$
and, for $x\in\{G,B\}$, $p^{(1)}(\,\cdot\,|x,(a_{1},a_{2}))=p^{(2)}(\,\cdot\,|x,(a_{1},a_{2}))=\delta_{x}$ (absorbing). Thus, under any $p\in\{p^{(1)},p^{(2)}\}$ the chain has two closed recurrent classes $\{G\}$ and $\{B\}$ (i.e., it is multi-chain).
Let the initial state be $S_{0}=s$. For stationary policies $\pi_{1}\in\Delta(\mathcal{A}_{1})$, $\pi_{2}\in\Delta(\mathcal{A}_{2})$, we specify them as
$\pi_{1}(1)=p\in[0,1],\qquad\pi_{2}(1)=q\in[0,1].$
Because $G,B$ are absorbing and $r_{1}\in\{0,1\}$ on $\{G,B\}$, the average reward of player 1 equals the probability (under the chosen kernel) of transitioning from $s$ to $G$ in one step. Thus we have that
$U(p,q)$ $\triangleq\min_{k\in\{1,2\}}\ \mathbb{P}_{p^{(k)}}\big(S_{1}=G\,\big|\,S_{0}=s,\,\pi_{1},\pi_{2}\big)$
$=\min\big\{\,p\,q+\varepsilon(1-p)(1-q),\ \varepsilon\,p\,q+(1-p)(1-q)\,\big\}.$
Then the robust average reward for player 1 is $U(p,q)$ and for player 2 is U(p,q)-U(p,q).
Lemma B.1.
For $U$ defined above and any $\varepsilon\in(0,1)$,
$\sup_{p\in[0,1]}\ \inf_{q\in[0,1]}U(p,q)\ =\ \frac{\varepsilon}{2},\qquad\inf_{q\in[0,1]}\ \sup_{p\in[0,1]}U(p,q)\ =\ \varepsilon.$
Proof.
Fix $p\in[0,1]$. For $q\in[0,1]$, $U(p,q)$ is the pointwise minimum of two affine functions of $q$, hence concave in $q$. Therefore the infimum over $q$ is attained at an endpoint:
$\inf_{q\in[0,1]}U(p,q)=\min\{\,U(p,0),\,U(p,1)\,\}=\min\{\varepsilon(1-p),\,\varepsilon p\}=\varepsilon\min\{p,1-p\}.$
Maximizing over $p$ gives
$\sup_{p\in[0,1]}\inf_{q\in[0,1]}U(p,q)=\sup_{p\in[0,1]}\varepsilon\min\{p,1-p\}=\varepsilon\cdot\frac{1}{2}=\frac{\varepsilon}{2}.$
Fix $q\in[0,1]$. As a function of $p$, $U(p,q)$ is again the pointwise minimum of two affine functions, hence concave and piecewise linear with a single kink at the solution of equality:
$p\,q+\varepsilon(1-p)(1-q)=\varepsilon p\,q+(1-p)(1-q)\ \Longleftrightarrow\ p=1-q.$
Thus
$\sup_{p\in[0,1]}U(p,q)=\max\Big\{\,U(0,q),\ U(1,q),\ U(1-q,q)\,\Big\}=\max\Big\{\,\varepsilon(1-q),\ \varepsilon q,\ (1+\varepsilon)q(1-q)\,\Big\}.$
Hence
$\inf_{q\in[0,1]}\ \sup_{p\in[0,1]}U(p,q)=\min_{q\in[0,1]}\ \max\Big\{\,\varepsilon(1-q),\ \varepsilon q,\ (1+\varepsilon)q(1-q)\,\Big\}.$
For $q\in\{0,1\}$ the maximum equals $\varepsilon$. For $q\in(0,1)$, $(1+\varepsilon)q(1-q)>\varepsilon\min\{q,1-q\}$ and thus the inner maximum is at least $\max\{\varepsilon q,\varepsilon(1-q)\}\geq\varepsilon/2$, while near $q=0$ or $1$ it is strictly larger than $\varepsilon/2$. The minimum over $q$ is attained at $q=0$ or $q=1$, and equals $\varepsilon$. ∎
Theorem B.1 (Nonexistence of a Nash equilibrium).
For the robust Markov game above, no (stationary mixed) Nash equilibrium exists.
Proof.
In a two-player zero-sum game, a stationary Nash equilibrium $(p^{\star},q^{\star})$ is a saddle point:
$U(p^{\star},q)\ \geq\ U(p^{\star},q^{\star})\ \geq\ U(p,q^{\star})\qquad\forall\,p,q\in[0,1],$
which implies
$\inf_{q}U(p^{\star},q)\ =\ U(p^{\star},q^{\star})\ =\ \sup_{p}U(p,q^{\star}).$
Consequently,
$\sup_{p}\inf_{q}U(p,q)\ \geq\ \inf_{q}U(p^{\star},q)\ =\ U(p^{\star},q^{\star})\ =\ \sup_{p}U(p,q^{\star})\ \geq\ \inf_{q}\sup_{p}U(p,q),$
so a saddle requires $\sup_{p}\inf_{q}U=\inf_{q}\sup_{p}U$. By Lemma B.1, $\sup_{p}\inf_{q}U=\varepsilon/2$ while $\inf_{q}\sup_{p}U=\varepsilon$, leading to a contradiction. Hence no saddle point exists and no Nash equilibrium exists. ∎
Appendix C Irreducible DR-MGs
Theorem C.1.
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the robust Bellman equation:
@@ Appendix C Irreducible DR-MGs
We denote $s^{\star}\triangleq\arg\min_{s\in\mathcal{S}}x(s)$, then it holds that
$x(s^{\star})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s).$ (30)
By our assumption, for any $s^{\prime}\in\mathcal{S}$, there exists some $k$ such that $((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}>0$. Then it holds that
$x(s^{\star})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s^{\star})=x(s^{\star}),$ (31)
@@
$x(s^{\star})=((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+(1-((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}})x(s^{\star}).$ (32)
@@
$((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\star})\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime}),$ (33)
which implies $x(s^{\star})=x(s^{\prime}),\forall s^{\prime}\in\mathcal{S}$, and thus $x=me$ for some constant $m$, i.e., $V^{\pi}_{\mathsf{P}_{w}}=me+V^{\pi}_{\mathsf{P}_{V}}$.
@@
Denote the optimal robust policy by $\pi^{\star}$, and the corresponding worst-case kernel by $\mathsf{P}^{\star}$. By Theorem 4.1, $(g^{\pi^{\star}}_{\mathcal{P}},h^{\pi^{\star}}_{\mathsf{P}^{\star}})$ is the solution to the robust Bellman equation of $\pi^{\star}$:
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)=r^{\pi^{\star}}(s)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{\pi^{\star}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})=\sum_{a}\pi^{\star}(a|s)\left(r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right).$ (41)
We now set $\pi^{\prime}(s)=\arg\max_{a}\{r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\},\forall s$, then we have that
$r^{\pi^{\prime}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})=g^{\pi^{\prime}}_{\mathcal{P}}\leq g^{\pi^{\star}}_{\mathcal{P}}$ $=\sum_{a}\pi^{\star}(a|s)\left(r(s,a)-h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right)$
$\leq r^{\pi^{\prime}}(s)-h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}}),$ (42)
@@
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)\leq\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})-\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})\leq\mathsf{P}^{\prime}(h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}).$ (43)
Now denote $x=h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}$, it holds that